a hollow conducting sphere of radius .12 meters has a uniform charge distributed on its surface. the electric field at a distance of .20 meters has a value of 10 n/c. how much charge is distributed on the sphere


1
Expert's answer
2021-04-12T06:58:27-0400
"E=k\frac{q}{r^2}\\10=9\cdot10^{9}\frac{q}{0.2^2}\\q=4.4\cdot10^{-11}C"